1/2(x+4)=1/3(3x-6)
Answer by checkley77(12569)

1/2(x+4)=1/3(3x-6)
(x+4)/2=(3x-6)/3
(x+4)/3=3(x-2)/3
(x+4)/2=x-2 cross multiply
2(x-2)=x+4
2x-4=x+4
2x-x=4+4
x=8 ans.
Proof:
(8+4)/2=(3*8-6)/3
12/2=(24-6)/3
6=18/3
6=6